World Company expects to operate at 60% of its productive capacity of 32,000 units per month. At this planned level, the company expects to use 12,000 standard hours of direct labor. Overhead is allocated to products using a predetermined standard rate of 0.625 direct labor hours per unit. At the 60% capacity level, the total budgeted cost includes $36,000 fixed overhead cost and $120,000 variable overhead cost. In the current month, the company incurred $116,000 actual overhead and 5,760 actual labor hours while producing 9,400 units. (Do not round intermediate calculations. Round "OH costs per DL hour" to 2 decimal places.)(1) Compute the predetermined standard overhead rate for total overhead. Predetermined OH rate Variable overhead costs Fixed
Answer #1

Answer:

Budget Production=32,000 units*60%

=32,000*60/100

=19,200 units

i) Given Variable overhead cost=$1,20,000

Fixed overhead cost =$ 36,000

Actual OH =$1,16,000

Now calculate Variable overhead rate and fixed overhead rate and Total overhead rate per unit Following:

Variable OH rate per unit=1,20,000/19,200=$6.25 per unit

Fixed OH rate per unit =36,000/19,200 =$1.875 per unit (or)     =(12,000 hours/19,200 units)*3=$1.875 per unit

Total OH cost per unit =$6.25+$1.875=$8.125 per unit

Fixed OH rate per hour =$36,000/12,000 h =$3 per labor hour

OVER HEAD VARIANCE

Particulars Standard OH actual production Actual OH Variance favourable/unfavourable
Variable OH =9,400 u*$6.25=$58,750 80,000(1,16,000-36,000) =80000-58750=21,250 Unfavourable
Fixed OH =5,760 h*$3=$17,280 36,000 =36000-17280=18,720 Unfavourable
Total OH Cost =$58,750+$17,280=$76,030 1,16,000 =21,250+18,720=39,970 Unfavourable